### v1

@@ 4.3. Knowledge Extraction from Conversation
CHOIR turns any Slack messages into a piece of reusable knowledge from conversation (DG1), allowing any members to participate in organizational memory maintenance(DG4). The follow-up discussion after a Q&A exchange is shared is likely to have answers to the question that the original questioner could not gain from CHOIR, which may be worth documenting.
When any lab member sees new knowledge worth keeping, they can summon CHOIR to request a documentation update by mentioning its handle, i.e., @CHOIR in a message ( C in Figure 5). At that point, CHOIR examines the previously shared Q&A together with the conversation up to the mention, identifies documentable content (filtering out personal or social chatter), and drafts a suggested update ( D in Figure 5). The requester can then edit this draft and press Suggest Update to forward it to the manager ( E in Figure 5). CHOIR can be summoned anywhere in Slack. For example, if a professor shared a writing tip as part of giving feedback to a student, a professor can summon CHOIR to document the knowledge, which is represented as a new start node in the bottom right corner of Figure 2. This feature simplifies the document update process, supporting DG1, by allowing any member to initiate updates and refine the knowledge rather than placing the entire responsibility on the manager. It lowers the barrier to documenting knowledge, thereby facilitating all lab members’ participation in organizational memory practices (DG4).
@@ 4.4. AI-assisted Document Update
This feature of AI-assisted update supports DG1 and DG4 in multiple ways. Once any member suggests an update from the Knowledge Extraction feature by pressing the “Suggest Update” button in Figure 5- E, CHOIR delivers it to the manager as a Slack DM message ( A in Figure 6) from CHOIR. The message presents the proposed update together with the contributor’s name and profile image, making the source and context explicit. At this stage, the manager can edit the suggestion, start the update process, or decline it, and these choices are visible not only in the DM but also as a message in the original channel so that other members are kept updated on how the update request is in progress. This final approval step, which the manager can only handle, ensures that only accurate information is documented, which is relevant to DG4.
Once the manager starts the update process by pressing the “Start Update Process” button, CHOIR will send follow-up messages one step at a time, guiding the manager to access which file to update, and how. First, the manager can select a target file from the repository or create a new file if the knowledge does not fit into an existing one ( B in Figure 6). If not selected, CHOIR will use the most relevant file depending on the similarity of the suggested update and document content. CHOIR then shows which file is selected and the proposed change in a diff-style message with inline emphasis on additions and removals. For each suggestion, the manager may edit the content directly, apply the change to the document, skip it to show the next relevant document — as there can be multiple places to update in a document, or stop the review entirely ( C in Figure 6). If the knowledge does not fit into an existing section, the interface provides an option to create a new section through a dedicated modal ( D in Figure 6). Applied updates—whether to an existing file, a new file, or a new section—are written back to the repository, and CHOIR posts a message to the original channel acknowledging that the update has been finished.
This feature implements DG1 by simplifying the document update process and automating many of the steps involved: (1) recommending where an update is needed, (2) proposing updated content and checking for conflicts with existing material, and (3) committing the changes at the point where new knowledge becomes available. In this way, managers can update documents directly within the flow of communication in Slack, without switching tools. While CHOIR automates much of the update process, it also preserves flexibility and managerial autonomy by allowing managers to edit the proposed content or make changes directly on GitHub. For example, if an update requires more than adding or revising text—such as changing headings or reorganizing the structure—CHOIR provides hyperlinks that open the file editor directly (“here” in D in Figure 6). Overall, this AI-assisted update feature enables managers to keep documents up to date with minimal effort while retaining full control over what becomes part of the organizational memory.
@@ 4.5. System Implementation
4.5.1. Architecture
CHOIR uses Slack as the front-end client where members interact with the system. The backend server is built with the JavaScript-based Slack Bolt [^2] https://api.slack.com/bolt framework on top of Express. In this setup, the server acts as an HTTP service that receives events from the Slack client as requests and returns messages as HTTP responses. The server is also connected to the lab’s GitHub repository through the GitHub REST API [^3] https://docs.github.com/en/rest, which serves as the canonical backend store for documentation. All documents in the repository are maintained in Markdown format, providing a lightweight and widely supported format. This stack allows Slack to serve as the entry point for everyday interactions, while GitHub provides persistent storage and version control for organizational memory.
4.5.2. Knowledge Retrieval and Generation
For Q&A, CHOIR maintains an embedding index over the documentation repository and performs retrieval-augmented generation conditioned on the user’s question and the retrieved text. Input handling is implemented through prompting: the LLM is instructed to decide whether an incoming message should be treated as a Q&A query or an update request. Responses to questions, knowledge extraction from conversations, and documentation update drafts are all generated via the OpenAI API. To protect privacy, user names are replaced with pseudonyms before any message is sent to the API. The exact prompts used for classification and generation are included in the Appendix for transparency and reproducibility.
@@ 5.1. Study Design
To explore these questions, we conducted an exploratory field deployment study that combined system usage logs with user interviews (Creswell and Clark, 2017), enabling us to capture both behavioral patterns and the reasoning behind user actions. We deployed CHOIR in active research laboratories to observe usage and experiences within participants’ real work environments (Rogers, 2011). This approach is particularly appropriate given our research focus on understanding user engagement and experience with a sociotechnical intervention in organizational contexts, where authentic workplace dynamics are crucial for meaningful insights.
@@ 5.2. Participants and Recruitment
We recruited university research lab directors through our social media and from formative study participants who expressed interest in the field study. Four lab directors initially agreed to participate; however, one later declined full deployment due to concerns about potential human-subject data leakage caused by human error. Upon confirmation from the director, we advertised the study in their slack worksapce and recruited the students from the lab, endorsed by the lab director. In total, four research laboratories participated in the user study, comprising 21 participants: three lab directors and 18 students. We refer to each unit and participant based on the labels as specified in Table 1. Since our study was conducted in the Summer of 2025, each lab had remote incoming graduate students joining during this period, allowing us to observe newcomer experiences.
All participants provided informed consent for interviews, system usage monitoring, and interaction log collection. The study was approved by our institution’s IRB. As compensation, lab directors received \$40 and students received \$20 upon completing the interviews.
@@ 5.3.1. Phase 1: Document Collection and Onboarding Sessions
Document Collection: Before the lab director onboarding sessions, we asked professors to share existing lab documents or materials they wished to integrate with CHOIR. We specifically requested only documents that could be made publicly available. D1 and D3 did not have any documents to start with, while D2 had a 1,650-word-long document that explained the lab director’s mentoring plan. To avoid a cold-start problem, we generated some documents based on information from each lab’s website and the graduate handbook that each university has, which included requirements, milestones, and timelines for graduate students. The research team converted these documents into markdown format, uploaded them to a public GitHub repository for each lab, and granted lab directors write access and students read access.
@@ 5.5. Data Analysis
Transcripts from post-deployment log-informed interviews were analyzed using reflexive thematic analysis (Braun and Clarke, 2019). Two researchers independently coded a subset of transcripts to establish themes, then collaboratively developed a coding framework applied to all interviews. Key themes focused on how participants used CHOIR for organizational memory, the barriers they faced, and their suggestions for improvement. We also compared interview themes with patterns in the system logs to understand not only what participants did, but also why they made particular choices and how they experienced the system.
@@ 6. Field Study Results
Overall, CHOIR successfully supported knowledge retrieval and updates to organizational memory during the study period, while also revealing nuanced challenges and opportunities for chatbot-mediated knowledge management systems, which we present below.
@@ 6.1. Q&A Interactions with CHOIR
During the one-month field study, participants submitted a total of 107 questions to CHOIR through both direct messages and public channels. Of these queries, 45 (42%) were successfully answered using the connected lab documentation, while CHOIR could not answer the remaining questions and revealed gaps in the existing knowledge base. Participants valued CHOIR’s effectiveness and reliability, with S2-1 noting that “it definitely improved the speed of retrieving the information compared to me reading through the document or searching online.” S4-2 emphasized the system’s accuracy: “if I asked a question and if it was there, you know, it would always give me a right answer or like summary of the documents.” D1 appreciated the quality of responses: “I think in general the answers were pretty good. When the response or when the like data was there, further details were there.
@@ 6.2.1. Q&A Sharing for Collecting Answers as well as Questions
Lab directors particularly appreciated the centralized Q&A channel for its organizational benefits. D2 valued the dedicated space for continuous interaction: “I like the Q&A channel. So it’s like those that’s for people to ask questions and I can answer them. I think I would like to keep like Q&A channel […] If people have questions, they can ask them there.” D3 emphasized the channel’s role in managing communication flow: “I think having this channel can kind of centralize all those messages in one place, so I like it.” While Q&A sharing was primarily designed for getting answers for unanswered questions, it supported DG3 in a way that we did not anticipate. Overall, deploying CHOIR in these labs encouraged and fostered the culture of sharing knowledge through Q&A using various channels: plan Slack communication, asking CHOIR in a public channel, and sharing the Q&A feature, successfully supporting our goal of facilitating interpersonal communication (DG3).
@@ 6.3.1. Perceived Value of Document Update via CHOIR
Participants recognized significant value in collaborative documentation, particularly appreciating the potential for preserving and sharing knowledge within research labs. The organizational value of maintaining documentation was widely acknowledged, with D1, who did not have any organizational memory beforehand, expressing satisfaction: “the documents. I’m very satisfied […] I think it’ll continue to live on and continue to be updated even after the study.” D2 emphasized the importance of capturing firsthand student experiences: “because they have firsthand information […] if they don’t document this, then the next person would have to figure it out by themselves again.” Students also appreciated the value of documentation, with S4-1 noting that “documentation is good. And another kind of documentation is all of our Slack messages. And it really sucks that those go away after three months or whatever. Because I feel like there’s actually a lot of organizational knowledge that is embedded within.
Directors particularly valued the AI-assisted update workflow, which reduced the effort required for documentation maintenance. D3 appreciated CHOIR’s editing suggestions: “CHOIR suggested me a sort of edit, which I liked […] I don’t have to spend a lot of time crafting.” D2 found the process straightforward, noting, “I just click add and I think it asks me to select a document.” All directors who either received update requests from students or initiated updates themselves successfully incorporated them into the lab documentation, demonstrating the effectiveness of the workflow when participants chose to engage.
@@ 6.3.3. Barriers to Student Contributions in Documentation
Although some students contributed to the knowledge when their lab director specifically requested, as in D2’s case above, students did not actively engage in initiating document updates; only three participants (S2-2, S2-3, S3-2) attempted to use the feature, all through private messages. These three members initiated a total of five updates during the study period. Although we expected students to request updates through the public Q&A channel, they primarily discovered documentation gaps during private Q&A with CHOIR, providing corrections based on their own knowledge before proposing updates to directors. For example, S3-2 noted, “I was asking about which courses to take, and CHOIR talked about the two required courses, but the course code was incorrect […] And so I just updated it.” This pattern validated the feasibility of CHOIR in enabling students to contribute to documentation.
@@ 7. Discussion
In this section, we reflect on our findings to understand the broader implications for designing AI-mediated organizational memory systems. Our field study revealed several key tensions and opportunities that extend beyond the specific technical features of CHOIR to more fundamental questions about how knowledge management systems can effectively support collaborative work in research organizations.
7.1. Privacy-Awareness Tension in Organizational Memory Systems
@@
Our findings revealed a fundamental tension between students’ privacy preferences and managers’ need for awareness in organizational memory systems. Because the majority of Q&A took place through direct messages to CHOIR without being shared, lab directors had limited visibility into knowledge gaps in the documentation. Directors expressed both their willingness and desire to update information that their students could not access through interactions with CHOIR. This challenge reflects a broader issue in collaborative work: maintaining awareness of individual and group activities is essential for successful collaboration (Dourish and Bellotti, 1992; Tam and Greenberg, 2006), yet our study shows that AI-mediated knowledge retrieval can suffer from the same limitation.
@@
This privacy-awareness tension stems from what communication researchers have identified as the fundamental trade-off between psychological safety and organizational transparency. Students’ behaviors—seeking private channels to avoid judgment and protecting others from perceived disruption—represent manifestations of psychological safety-seeking in hierarchical organizations. Research has consistently demonstrated that anonymous communication reduces social pressure and enables more honest self-expression, particularly when interacting with authority figures (Suler, 2004; Clark-Gordon et al., 2019). However, this creates what organizational behavior scholars have identified as a fundamental paradox: while privacy enables honest participation, complete transparency can also discourage participation by making members feel overly monitored (Bernstein, 2012).
@@
These findings suggest that future organizational memory systems must balance individual privacy with the members’ benefit coming from reinforced organizational memory. Our participants suggested privacy-preserving awareness mechanisms, one of which is to periodically report questions at high levels, which would inform managers about question patterns without compromising individual privacy. Given the students’ preference for direction communication over CHOIR-mediated communication, this alternative can be more effective for lab directors to be aware of the gaps in organizational memory. This approach preserves student autonomy while providing managers with actionable insights about knowledge gaps and documentation needs.

### v2

@@ 4.3. Knowledge Extraction from Conversation
CHOIR supports turning any Slack messages into reusable knowledge (DG1), allowing members to participate in organizational memory maintenance (DG4). Follow-up discussions after Q&A sharing often contain answers worth documenting. When any lab member sees new knowledge worth keeping, they can summon CHOIR by mentioning @CHOIR ( C in Figure 5). CHOIR then examines the conversation, identifies documentable content, and drafts a suggested update ( D). The requester can edit this draft and press Suggest Update to forward it to the manager ( E). CHOIR can be summoned anywhere in Slack—for example, if a professor shares a writing tip while giving feedback to a student, the professor can summon CHOIR to document that knowledge, as represented by the new start node in the bottom right corner of Figure 2.
@@ 4.4. AI-assisted Document Update
This feature of AI-assisted update supports DG1 and DG4 in multiple ways. Once any member suggests an update from the Knowledge Extraction feature by pressing the “Suggest Update” button ( E in Figure 5), CHOIR delivers it to the manager as a Slack DM message ( A in Figure 6) from CHOIR. The message presents the proposed update together with the contributor’s name and profile image, making the source and context explicit. At this stage, the manager can edit the suggestion, start the update process, or decline it, and these choices are visible not only in the DM but also as a message in the original channel so that other members are kept updated on how the update request is in progress. This final approval step, which the manager can only handle, is designed to help maintain the accuracy of documented information, which is relevant to DG4.
Once the manager starts the update process by pressing the “Start Update Process” button, CHOIR will send follow-up messages one step at a time, guiding the manager to access which file to update, and how. First, the manager can select a target file from the repository or create a new file if the knowledge does not fit into an existing one ( B in Figure 6). If not selected, CHOIR will use the most relevant file depending on the similarity of the suggested update and document content. CHOIR then shows which file is selected and the proposed change in a diff-style message with inline emphasis on additions and removals. For each suggestion, the manager may edit the content directly, apply the change to the document, skip it to show the next relevant document — as there can be multiple places to update in a document, or stop the review entirely ( C). If the knowledge does not fit into an existing section, the interface provides an option to create a new section through a dedicated modal ( D). Applied updates—whether to an existing file, a new file, or a new section—are written back to the repository, and CHOIR posts a message to the original channel acknowledging that the update has been finished.
This feature implements DG1 by automating key steps: recommending where updates are needed, proposing content, and committing changes. Managers can update documents directly in Slack without switching tools. CHOIR preserves managerial autonomy by allowing edits to proposed content or, for more complex tasks such as changing headings or reorganizing structure, enabling direct changes on GitHub via provided hyperlinks (“here” in D in Figure 6).
@@ 4.5. System Implementation
CHOIR uses Slack as the front-end client and is built with the Slack Bolt [^2] https://api.slack.com/bolt framework. The backend connects to the lab’s GitHub repository through the GitHub REST API [^3] https://docs.github.com/en/rest, which serves as the persistent store for documentation in Markdown format. This stack allows Slack to serve as the entry point for everyday interactions, while GitHub provides version control for organizational memory.
For Q&A, CHOIR maintains an embedding index over the documentation repository and performs retrieval-augmented generation conditioned on the user’s question and the retrieved text. We use GPT-4o (temperature=0, max_tokens=1000) for text generation and text-embedding-3-small for document embeddings via the OpenAI API. Input handling is implemented through prompting: the LLM is instructed to decide whether an incoming message should be treated as a Q&A query or an update request. Responses to questions, knowledge extraction from conversations, and documentation update drafts are all generated via the OpenAI API. To protect privacy, user names are replaced with pseudonyms before any message is sent to the API. The complete system prompts are provided in Appendix A.
@@ 5.1. Study Design
To explore these questions, we conducted an exploratory field deployment study that combined analysis of system usage logs with user interviews (Creswell and Clark, 2017), enabling us to capture both behavioral patterns and the reasoning behind user actions. We deployed CHOIR in active research laboratories to observe usage and experiences within participants’ real work environments (Rogers, 2011). This field study approach reflects the often underestimated difficulty of evaluating CSCW applications: organizational systems must be assessed within authentic work contexts, as in-lab experiments that attempt to simulate ecologically valid environments cannot reproduce the social dynamics and group interactions that unfold over weeks and months (Grudin, 1988).
Our evaluation focuses on understanding how CHOIR mediates organizational knowledge management practices, thereby demonstrating the feasibility of CHOIR and the benefits observed through user behaviors and interview responses. Field studies of organizational memory systems have long shown how such systems augment collective memory within real organizational contexts and workflows (Ackerman, 1998). While field deployments cannot establish causal relationships due to the challenges of creating controlled baselines, they enable a holistic understanding via a descriptive approach (Lazar et al., 2017) that characterizes how CHOIR mediates organizational communication and memory practices in situ. Many social computing studies that involve a group of users, an organization, or a community, similarly choose to evaluate systems through deploying the system in the wild and takes the descriptive approach (Mahar et al., 2018; Munson, 2008; Jasim et al., 2021; Bali et al., 2023; Mahyar et al., 2018; Kriplean et al., 2012; Cranshaw et al., 2017). In studying sociotechnical interventions in organizational contexts, authentic workplace dynamics provide crucial insights into user engagement and experience with collaborative knowledge management systems.
@@ 5.2. Participants and Recruitment
We recruited university research lab directors through our social media and from formative study participants who expressed interest in the field study. Four research laboratories participated in the user study, comprising 22 participants: four lab directors and 18 students. Upon confirmation from each director, we advertised the study in their Slack workspace and recruited students from the lab, endorsed by the lab director. We refer to each unit and participant based on the labels as specified in Table 1. The one-month deployment was conducted during the Summer of 2025, a transitional academic period when each lab had remote incoming graduate students joining. This timing allowed us to observe how CHOIR supports early-stage onboarding and knowledge exchange when new members have many initial questions.
All participants provided informed consent for interviews, system usage monitoring, and interaction log collection. The study was approved by our institution’s IRB. As compensation, lab directors received \$40 and students received \$20 upon completing the interviews. One participating lab (L4) was led by the corresponding author, who served as the lab director. To mitigate any potential pressure that L4 participants might feel due to power dynamics or a desire to be favorable toward their peers, the director limited their role to answering questions and making changes to documents and did not otherwise intervene in the study procedures. Similarly, the first and second authors, who are members of L4, did not use CHOIR in this field deployment at all. All L4 participants voluntarily consented to the study independently through the standard procedure of advertisement, recruitment, and onboarding sessions, moderated by the first and second student authors. We further discuss potential limitations of this setup in the Discussion section.
@@ 5.3.1. Phase 1: Document Collection and Onboarding Sessions
Document Collection: Before the lab director onboarding sessions, we asked professors to share existing lab documents or materials they wished to integrate with CHOIR. We specifically requested only documents that could be made publicly available. The participating labs exhibited diverse documentation practices. D1 and D3 had no existing organizational memory systems and primarily relied on informal communication channels such as Slack messages, verbal exchanges, or external resources (e.g., department handbooks). To address this and avoid a cold-start problem, the research team worked with them to create initial seed documents based on available resources (e.g., lab websites and departmental handbooks), ensuring that CHOIR would contain useful content immediately upon deployment. In contrast, D2 and D4 already maintained web-based documentation, including a lab webpage and a GitHub repository containing the director’s advice, mentoring plan, and lab policies. However, during the study period, they voluntarily used CHOIR rather than their existing systems. The research team, with the directors’ approval and guidance, converted all provided materials into markdown files, uploaded them to a public GitHub repository for each lab, and granted lab directors write access and students read access.
@@ 5.5. Data Analysis
Transcripts from post-deployment log-informed interviews were analyzed using reflexive thematic analysis (Braun and Clarke, 2019). Two researchers independently coded a subset of transcripts to establish themes, then collaboratively developed a coding framework applied to all interviews. Key themes focused on how participants used CHOIR for organizational memory, the barriers they faced, and their suggestions for improvement.
@@ 6. Field Study Results
Overall, CHOIR supported knowledge retrieval and updates to organizational memory during the study period, while also revealing nuanced challenges and opportunities for chatbot-mediated knowledge management systems, which we present below.
@@ 6.1. Q&A Interactions with CHOIR
During the one-month field study, participants submitted a total of 107 questions to CHOIR through both direct messages and public channels. Of these queries, 45 (42%) were answered using the connected lab documentation, while CHOIR could not answer the remaining questions and revealed gaps in the existing knowledge base. Participants valued CHOIR’s effectiveness and reliability, with S2-1 noting that “it definitely improved the speed of retrieving the information compared to me reading through the document or searching online.” S1-2 appreciated its accuracy with reference: “if the information is already in the handbook, it’s pretty straightforward to get those information, and I like that it provided a reference […] That helped me to get the accurate part information.” D1 appreciated the quality of responses: “I think in general the answers were pretty good. When the response or when the like data was there, further details were there.
@@ 6.2.1. Q&A Sharing for Collecting Answers as well as Questions
Lab directors particularly appreciated the centralized Q&A channel for its organizational benefits. D2 valued the dedicated space for continuous interaction: “I like the Q&A channel. So it’s like those that’s for people to ask questions and I can answer them. I think I would like to keep like Q&A channel […] If people have questions, they can ask them there.” D3 emphasized the channel’s role in managing communication flow: “I think having this channel can kind of centralize all those messages in one place, so I like it.” While Q&A sharing was primarily designed for getting answers for unanswered questions, it supported DG3 in a way that we did not anticipate. Overall, deploying CHOIR in these labs encouraged and fostered the culture of sharing knowledge through Q&A using various channels: plan Slack communication, asking CHOIR in a public channel, and sharing the Q&A feature, supporting our goal of facilitating interpersonal communication (DG3).
@@ 6.3.1. Perceived Value of Document Update via CHOIR
Participants recognized significant value in collaborative documentation, particularly appreciating the potential for preserving and sharing knowledge within research labs. The organizational value of maintaining documentation was widely acknowledged, with D1, who did not have any organizational memory beforehand, expressing satisfaction: “the documents. I’m very satisfied […] I think it’ll continue to live on and continue to be updated even after the study.” D2 emphasized the importance of capturing firsthand student experiences: “because they have firsthand information […] if they don’t document this, then the next person would have to figure it out by themselves again.” S1-1 highlighted the cumulative value for future members: “the more questions asked and updated, the more useful it becomes for future students.
Directors particularly valued the AI-assisted update workflow, which reduced the effort required for documentation maintenance. D3 appreciated CHOIR’s editing suggestions: “CHOIR suggested me a sort of edit, which I liked […] I don’t have to spend a lot of time crafting.” D2 found the process straightforward, noting, “I just click add and I think it asks me to select a document.” Directors who either received update requests from students or initiated updates themselves incorporated them into the lab documentation, suggesting the effectiveness of the workflow when participants chose to engage.
@@ 6.3.3. Barriers to Student Contributions in Documentation
Although some students contributed to the knowledge when their lab director specifically requested, as in D2’s case above, students did not actively engage in initiating document updates; only three participants (S2-2, S2-3, S3-2) attempted to use the feature, all through private messages. These three members initiated a total of five updates during the study period. Although we expected students to request updates through the public Q&A channel, they primarily discovered documentation gaps during private Q&A with CHOIR, providing corrections based on their own knowledge before proposing updates to directors. For example, S3-2 noted, “I was asking about which courses to take, and CHOIR talked about the two required courses, but the course code was incorrect […] And so I just updated it.” This pattern suggests the feasibility of CHOIR in enabling students to contribute to documentation.
@@ 7. Discussion
We reflect on our findings to understand implications for designing AI-mediated organizational memory systems. Our field study revealed several key tensions and opportunities that extend beyond CHOIR’s technical features to more fundamental questions about how knowledge management systems can support collaborative work in research organizations.
7.1. Privacy-Awareness Tension in AI-mediated Organizational Memory
@@
Our findings revealed a fundamental tension between students’ privacy preferences and managers’ need for awareness in organizational memory systems. Because the majority of Q&A took place through direct messages to CHOIR without being shared, lab directors had limited visibility into knowledge gaps in the documentation. Directors expressed willingness to update information that their students could not access through interactions with CHOIR. This challenge reflects a broader issue in collaborative work: maintaining awareness of individual and group activities is essential for successful collaboration (Dourish and Bellotti, 1992; Tam and Greenberg, 2006), yet our study shows that AI-mediated knowledge retrieval can suffer from the same limitation.
Our analysis suggests that AI intermediaries such as CHOIR reconfigure organizational memory (OM) practices by redistributing visibility across members of an organization. By enabling workers to pose questions directly to an AI agent, these systems increase the visibility of existing organizational knowledge to questioners while reducing the workload of searching or manually locating information. However, this enhanced accessibility simultaneously diminishes members’ organic awareness of knowledge gaps, which was previously supported through in-person interactions. This redistribution produces new asymmetries and hidden dynamics in a group, requiring members to negotiate trade-offs between individual privacy, exposure of information needs, and the collective benefits of shared organizational knowledge. Further, the interaction traces generated through AI-mediated OM—such as question-asking histories—can serve as a valuable resource for identifying missing knowledge or understanding evolving information needs within the organization. Yet leveraging such data also risks inadvertently creating forms of workplace surveillance, particularly when anonymity is contained in a smaller group. These dynamics reveal that AI-mediated OM systems do not merely support knowledge access; they actively reshape the visibility, accountability, and communication of knowledge sharing practice, offering a lens for understanding how algorithmic intermediaries influence sociotechnical dimensions of OM.
@@
This redistribution of visibility intersects with a long-standing tension in communication research between psychological safety and organizational transparency. The tension stems from what communication researchers have identified as the fundamental trade-off between psychological safety and organizational transparency. Students’ behaviors—seeking private channels to avoid judgment and protecting others from perceived disruption—represent manifestations of psychological safety-seeking in hierarchical organizations. Research has consistently demonstrated that anonymous communication reduces social pressure and enables more honest self-expression, particularly when interacting with authority figures (Suler, 2004; Clark-Gordon et al., 2019). However, this creates what organizational behavior scholars have identified as a fundamental paradox: while privacy enables honest participation, complete transparency can also discourage participation by making members feel overly monitored (Bernstein, 2012).
@@
These findings suggest that future organizational memory systems must balance individual privacy with the members’ benefit coming from reinforced organizational memory. Our participants suggested privacy-preserving awareness mechanisms, one of which is to periodically report questions at high levels, which would inform managers about question patterns without compromising individual privacy. Given the students’ preference for direction communication over CHOIR-mediated communication, this alternative could be more effective for lab directors to be aware of the gaps in organizational memory. This approach preserves student autonomy while providing managers with actionable insights about knowledge gaps and documentation needs.